# Global decomposition of networks into multiple cores formed by local hubs

### v2

@@ IV Summary and outlook
We have proposed a method to uncover the core-periphery structure of networks through network decomposition centered around local hubs. Compared with other edge or node centralities such as edge betweenness and degree-product, our hub-centrality-product rule ensures the existence of a series of natural cutoffs in the form of a cusp on the giant component size versus the fraction of removed edges. The cusp point, unambiguously defined as the transition point of removed edges with zero versus nonzero hub-centrality product, signals the breaking point of a shell from a backbone. Our local-edge decomposition method repeats this process for each backbone as a brand-new network until we run out of edges to be removed. We have demonstrated the properties and implications of the method with a collaboration network as a representative example, in particular, compared with the celebrated k-core decomposition.
Naturally, the decomposition yields the core-periphery separation and we have introduced a principled way to pinpoint the boundary by the core-periphery score accounting the similarity to the ideal core-periphery structure with respect to the background edge density; in particular, we would like to emphasize again that our local-edge decomposition shines when it is combined with community division, by preparing the groups with appropriate sizes to handle with the decomposition in advance. We believe that the combination leads to the most natural way to extract so-called multiple core-periphery structures [19, 20, 21, 22], and the extension to the coarse-grained communities as supernodes clearly demonstrates the intermingled structure of core-periphery and communities [17]. From the result, we have clearly shown the merit of embracing local hubs in structurally dissecting networks, as in the previously reported effect on dynamical properties [23, 7]. Some related previous works by others include Ref. [38] on community structures, where so-called “whiskers” can be recognized as part of the periphery despite having relatively high degrees because it is based on relative connections. This approach may align in many aspects with the study that utilizes the hub centrality. Most of all, we hope that this type of perspective regarding locally important substructures, such as the concept of hidden dependency between nodes from it [62], gets more attention from the network science community.
@@ Acknowledgments
This work was supported by the National Research Foundation (NRF) of Korea under Grant Nos. NRF-2021R1C1C1004132 (S.H.L.) and NRF-2022R1A4A1030660 (W.J. and S.H.L.). The authors thank Ludvig Lizana for the discussion regarding the concept of using supernodes from network communities in a coarse-grained level, and Heetae Kim for insightful comments including correcting typos.

### v3

@@ IV Summary and Outlook
The most important insight from our analysis is the possibility of using the concept of local hubs, defined with respect to the degrees of the neighbors of a node of interest. We believe that the local hubs can be seeds (potentially corresponding to the local cores in terms of core-periphery separation) for a different type of network decomposition, compared with conventional ones such as the $k$-core decomposition from the global perspective. For that purpose, we have proposed a method to uncover the core-periphery structure of networks through network decomposition centered around local hubs. Compared with other edge or node centralities such as edge betweenness and degree-product, our hub-centrality-product rule ensures the existence of a series of natural cutoffs in the form of a cusp on the giant component size versus the fraction of removed edges. The cusp point, unambiguously defined as the transition point of removed edges with zero versus nonzero hub-centrality product, signals the breaking point of a shell from a backbone. Our local-edge decomposition method repeats this process for each backbone as a brand-new network until we run out of edges to be removed. We have demonstrated the properties and implications of the method with a collaboration network as a representative example, in particular, compared with the celebrated k-core decomposition.
Naturally, the decomposition yields the core-periphery separation and we have introduced a principled way to pinpoint the boundary by the core-periphery score accounting the similarity to the ideal core-periphery structure with respect to the background edge density; in particular, we emphasize again that our local-edge decomposition shines when it is combined with community division, by preparing the groups with appropriate sizes to handle with the decomposition in advance. We believe that the combination leads to the most natural way to extract so-called multiple core-periphery structures [19, 20, 21, 22], and the extension to the coarse-grained communities as supernodes clearly demonstrates the intermingled structure of core-periphery and communities [17]. From the result, we have clearly shown the merit of embracing local hubs in structurally dissecting networks, as in the previously reported effect on dynamical properties [23, 7]. Some related previous works by others include Ref. [38] on community structures, where so-called “whiskers” can be recognized as part of the periphery despite having relatively high degrees because it is based on relative connections. This approach may align in many aspects with the study that utilizes the hub centrality. Moreover, a comprehensive comparison between our approaches (LED and its application to detect core-periphery structures) and existing methods [14, 15, 16, 17, 18, 19, 20, 21, 22] will provide valuable insights in the core-periphery structure, and we leave it as foreseeable and plausible future research. Most of all, we hope that this type of perspective regarding locally important substructures, such as the concept of hidden dependency between nodes from it [62], gets more attention from the network science community.
@@ Acknowledgments
This work was supported by the National Research Foundation (NRF) of Korea under Grant No. RS-2021-NR061247. The authors thank Ludvig Lizana for the discussion regarding the concept of using supernodes from network communities in a coarse-grained level, Heetae Kim for insightful comments including correcting typos, and an anonymous referee for valuable suggestions for future work.